Solve.
A garden area is 30 feet long and 20 feet wide. A path of uniform width is set around the edge. If the remaining garden area is 400 feet squared, what is the width of the path?''.

Answer by scott8148(6628)

let x="width of path" __ (30-2x)(20-2x)=400 __ 600-100x+4x^2=400

x^2-25x+50=0

use quadratic formula to find x